# L’approximation arithmétique de Fujita sur courbes adéliques

### 27 Jan 2024 v1

@@ 4.2 Fujita approximation of graded linear series
Let $p\in\mathbb{N}$ such that $E_{\operatorname{HN},p}^{0}\neq 0$, and $\overline{E}_{\bullet}^{(p)}=\bigoplus\limits_{n\geqslant 0}\overline{E}^{(p)}_{\operatorname{HN},np}$ be the graded $K$-algebra generated by $E_{\operatorname{HN},p}^{0}$, which is considered as a normed graded linear series over $S_{0}$. At the same time, it is also able to be considered as a normed graded linear series over a general proper adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ equipped with the subspace norms of $\bigoplus\limits_{D\geqslant 0}\overline{E}_{D}$. In the both senses, they have the same arithmetic volume by Proposition 3.4.3 if $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$.
The following result is an analogue of [2, Theorem 4.1] over $S_{0}$, whose proof is quite similar to the ancient one.
@@
$\widehat{\operatorname{vol}}(\overline{L})=\sup_{p\in\mathbb{N}}\widehat{\operatorname{vol}}(\overline{E}_{\bullet}^{(p)}),$
where $\overline{E}_{\bullet}^{(p)}$ is defined as above, which is a normed graded linear series over $S_{0}$.
@@
In order to prove the assertion, we will prove that the restriction of $\nu_{\mathrm{sp}}$ on $[0,+\infty)$ is able to be approximated well by $\nu_{\mathrm{sp}}^{(p)}$.
@@
be a division of the interval $[0,\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L}))$ such that
@@
where the last equality is from [1, Chap. IV §5 $n^{\circ}$ 12 Proposition 22].
@@
$\max\{t,0\}-\max\{t-\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L}),0\}$
when $j\to\infty$. By definition, the support of $\nu_{\mathrm{sp}}$ is bounded from above by $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})$, we obtain
@@
For a general graded linear series, if it satisfies the property of Proposition 4.2, we say that this graded linear series satisfies Fujita approximation. We would like to emphasize that this notation is different from those in [5, Definition 3.7] and [6, §6].
@@ 5 Fujita approximation of adelic line bundles
From this Fujita approximation of graded linear series proved in Proposition 4.2, we will prove the arithmetic Fujita approximation of an arithmetically big line bundle.
@@ 5.1 Fujita approximation of big adelic line bundles
In this part, we will prove the arithmetic Fujita approximation of usual adelic line bundles. First, we prove a nef version.
@@
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. Let $X$ be a projective variety with $\dim(X)=d$, and $\overline{L}$ be an arithmetically big line bundle on $X$ over $S$. Then for every $\epsilon>0$, there exists a birational morphism $v:X^{\prime}\rightarrow X$ together with a positive integer $p$, such that the decomposition $v^{*}(\overline{L}^{\otimes p})=\overline{M}\otimes\overline{G}$ satisfies:
@@
We devide the proof into three parts.
@@
$X_{p}=\operatorname{Proj}\left(\operatorname{Im}\left(\bigoplus_{n\geqslant 0}\pi^{*}(E_{\operatorname{HN},np}^{p})\longrightarrow\bigoplus_{n\geqslant 0}L^{\otimes np}\right)\right).$
@@
For any $n\geqslant 1$ and $w\in\Omega$, let .$\|\raisebox{1.72218pt}{.}\|_{n,w}$ be the quotient norm on $G_{p,w}=G_{p}\times_{\operatorname{Spec}K}\operatorname{Spec}K_{w}$ induced by the surjective homomorphism
@@
In fact, for any $x\in X_{p,w}(\overline{K}_{w})$ outside the exceptional divisor, it corresponds to a $1$-dimensional quotient of $E_{\operatorname{HN},p}^{0}$. For any $n\geqslant 1$, this quotient induces a $1$-dimensional quotient $\ell_{n,x}$ of $E_{\operatorname{HN},np}^{(p)}$. By Hahn-Banach theorem, every linear form on $\ker(E_{\operatorname{HN},np}^{(p)}\rightarrow\ell_{n,x})$ can be extended to a linear form on $E_{\operatorname{HN},np}^{(p)}$. In other words, there exists a section $w\in E_{\operatorname{HN},np}^{(p)}$ whose image in $G_{p,w}^{\otimes n}(x)$ has norm $\|w\|_{\sup,w}$. As a section of $L_{w}^{\otimes np}$ on $X_{v}(\overline{K}_{w})$, we have $\|w(x)\|_{w}\leqslant\|w\|_{\sup,w}$, where .$\|\raisebox{1.72218pt}{.}\|_{w}$ is the metric on $\overline{L}$ indexed by $v\in\Omega$. Then for any section $u\in G_{p,w}$ over a neighborhood of $x$, we may assume that $u(x)^{\otimes n}$ equals to the image of $w(x)$ in $G_{p,w}^{\otimes n}(x)$ by dilation, then we have
@@
Let $\alpha_{n}=(\|\raisebox{1.72218pt}{.}\|_{n,w})_{W\in\Omega}$ be an adelic metric on $G_{p}$, and we define $(M_{p},\beta_{n})=\phi_{p}^{*}(\overline{L})\otimes(G_{p},\alpha_{n})^{\vee}$. In this sense, $s$ is an effective section in $(M_{p},\beta_{n})$, then it is an effective line bundle.
@@
$\widehat{\operatorname{vol}}(E_{\bullet}^{(p)},\alpha_{n})=\lim_{m\rightarrow\infty}\frac{\widehat{\deg}_{+}(E_{\operatorname{HN},mp}^{(p)},(\|\raisebox{1.72218pt}{.}\|_{n,w,\sup})_{w\in\Omega})}{m^{d+1}/(d+1)!},$
and .$\|\raisebox{1.72218pt}{.}\|_{n,w,\sup}$ is the superior norm induced by $\alpha_{n}=(\|\raisebox{1.72218pt}{.}\|_{n,w})_{w\in\Omega}$ on $G_{p}$. Then by [2, Corollary A.2], we have
@@
for infinitely many $n\geqslant N_{0}$. In addition, by Proposition 3.4.4 we obtain that $(G_{p},\alpha_{n})$ is arithmetically nef, since $\widehat{\mu}_{\min}(\overline{E}_{\operatorname{HN},np}^{(p)})\geqslant 0$ for all $n\geqslant 1$, then we deduce the positive asymptotic minimal slope.
@@
Next, we obtain the following ample version from Theorem 5.1 directly.
@@
Let $\overline{B}$ be an arithmetically big line bundle on $X$. Then by [7, Corollary 6.4.10], for $m\in\mathbb{N}$ sufficiently large, we have $\overline{L}^{\otimes m}\otimes\overline{B}^{\vee}$ is arithmetically big. By Theorem 5.1, for any $\epsilon>0$, there exists a birational morphism $\phi:X^{\prime}\rightarrow X$, a $p\in\mathbb{N}$, and a decomposition
@@
By the continuity of the arithmetic volume function (see [7, Corollary 6.4.10]), we have the result. ∎
@@ 5.2 Case of $\mathbb{K}$-line bundles
In this part we will prove the arithmetic Fujita approximation is verified for adelic $\mathbb{Q}$-line bundles and $\mathbb{R}$-line bundles over adelic curve.
@@
First, we prove the case of adelic $\mathbb{Q}$-line bundle, which is a quite straightforward from the usual case.
@@
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. Let $X$ be a projective variety over $\operatorname{Spec}K$ with $\dim(X)=d$, and $\overline{L}$ be an arithmetically big adelic $\mathbb{Q}$-line bundle on $X$ over $S$. For every $\epsilon>0$, there exists a birational morphism $\phi:X^{\prime}\rightarrow X$ together with a $p\in\mathbb{N}^{+}$, such that the decomposition $\phi^{*}(\overline{L}^{\otimes p})=\overline{M}\otimes\overline{G}$ satisfies:
@@
    $\overline{M}$ and $\overline{G}$ are adelic $\mathbb{Q}$-line bundles, where $\overline{M}$ is effective and $\overline{G}$ is arithmetically ample;
@@
By definition, there exists an $m\in\mathbb{N}$, such that $\overline{L}^{\otimes m}$ is a usual arithmetically big adelic line bundle. Then we obtain the assertion by applying Corollary 5.1 to $\overline{L}^{\otimes m}$. ∎
@@
Next, we prove the case of adelic $\mathbb{R}$-line bundle, and see [7, §6.4.3] for an introduction to this notion. For this target, we prove an auxiliary lemma first, which allows us to approximate an arithmetically big $\mathbb{R}$-line bundle by an arithmetically big $\mathbb{Q}$-line bundle along an effective direction.
@@
The adelic $\mathbb{R}$-line bundle $\overline{L}$ is arithmetically big, so $L$ is a big $\mathbb{R}$-line bundle. By an equivalent definition [14, Definition 2.2.21], we have $L=\sum\limits_{i=1}^{r}a_{i}L_{i}$, where $a_{1},\ldots,a_{r}>0$ and $L_{1},\ldots,L_{r}$ are usual big line bundles. By Kodaira’s decomposition [14, Corollary 2.2.7], there exists an $m\in\mathbb{N}$, such that for all $i=1,\ldots,r$, we have $mL_{i}=A_{i}+M_{i}$, where $A_{i}$ is ample and $M_{i}$ is effective.
We equip metrics on $A_{1},\ldots,A_{r},M_{1},\ldots,M_{r}$ over $S$, such that $\overline{A}_{1},\ldots,\overline{A}_{r}$ are arithmetically ample, and $\overline{M}_{1},\ldots,\overline{M}_{r}$ are arithmetically effective. Then for any $b_{1},\ldots,b_{r},c_{1},\ldots,c_{r}\in\mathbb{R}^{+}$, we have $b_{1}\overline{A}_{1}+\cdots+b_{r}\overline{A}_{r}+c_{1}\overline{M}_{1}+\cdots+c_{r}\overline{M}_{r}$ is effective.
By the continuity of $\widehat{\operatorname{vol}}(\raisebox{1.72218pt}{.})$ from [7, Corollary 6.4.24], for all $\epsilon>0$, there exist $\delta_{1},\ldots,\delta_{r},\eta_{1},\ldots,\eta_{r}>0$, such that for all $t_{i}\in(0,\delta_{i})$, $s_{i}\in(0,\eta_{i})$ for $i=1,\ldots,r$, we always have
$0<\widehat{\operatorname{vol}}(m\overline{L})-\widehat{\operatorname{vol}}\left(m\overline{L}-\sum_{i=1}^{r}t_{i}\overline{A}_{i}-\sum_{i=1}^{r}s_{i}\overline{M}_{i}\right)<\epsilon.$
If there does not exist an arithmetically big $\mathbb{Q}$-line bundle satisfying the requirements of the lemma, then $m\overline{L}-\sum\limits_{i=1}^{r}t_{i}\overline{A}_{i}-\sum\limits_{i=1}^{r}s_{i}\overline{M}_{i}$ can never be a $\mathbb{Q}$-line bundle for all $t_{i}\in(0,\delta_{i})$ and $s_{i}\in(0,\eta_{i})$ $i=1,\ldots,r$. It means there is no $\mathbb{Q}$-point in the region
$(ma_{1}-\delta_{1},ma_{1})\times\cdots\times(ma_{r}-\delta_{r},ma_{r})\times(ma_{1}-\eta_{1},ma_{1})\times\cdots\times(ma_{r}-\eta_{r},ma_{r})\subset\mathbb{R}^{2r},$
which has a strict positive Lebesgue measure. Then it contradicts to the fact that $\mathbb{Q}^{2r}$ is dense in $\mathbb{R}^{2r}$ with respect to the Euclidean topology.
By choosing the constant $\epsilon$ sufficiently small, we pick a family $t_{1},\ldots,t_{r},s_{1},\ldots,s_{r}\in\mathbb{R}^{+}$ such that $m\overline{L}-\sum\limits_{i=1}^{r}t_{i}\overline{A}_{i}-\sum\limits_{i=1}^{r}s_{i}\overline{M}_{i}$ is an arithmetically big $\mathbb{Q}$-line bundle. By the above construction, $\sum\limits_{i=1}^{r}t_{i}\overline{A}_{i}+\sum\limits_{i=1}^{r}s_{i}\overline{M}_{i}$ is effective. Let
$\overline{L}^{\prime}=\overline{L}-\frac{1}{m}\sum\limits_{i=1}^{r}t_{i}\overline{A}_{i}-\frac{1}{m}\sum\limits_{i=1}^{r}s_{i}\overline{M}_{i},$
@@
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. Let $X$ be a projective variety over $\operatorname{Spec}K$ with $\dim(X)=d$, and $\overline{L}$ be an arithmetically big adelic $\mathbb{R}$-line bundle on $X$ over $S$. For every $\epsilon>0$, there exists a birational morphism $\phi:X^{\prime}\rightarrow X$ together with a $p\in\mathbb{N}^{+}$, such that the decomposition $\phi^{*}(\overline{L}^{\otimes p})=\overline{M}\otimes\overline{G}$ satisfies:
@@
    $\overline{M}$ and $\overline{G}$ are adelic $\mathbb{R}$-line bundles, where $\overline{M}$ is effective and $\overline{G}$ is arithmetically ample;

### 17 Mar 2024 v2

@@ 4.2 Fujita approximation of graded linear series
Let $p\in\mathbb{N}$ such that $E_{\operatorname{HN},p}^{0}\neq 0$, where we follow the notation in (9). In this case, we define $\overline{E}_{\bullet}^{(p)}=\bigoplus\limits_{n\geqslant 0}\overline{E}^{(p)}_{\operatorname{HN},np}$ to be the graded $K$-algebra generated by $E_{\operatorname{HN},p}^{0}$, which is considered as a normed graded linear series over $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ defined in Example 2.1.1.
The following result is an analogue of [4, Theorem 4.1] over $S_{0}$, whose proof is quite similar to the ancient one.
@@
$\widehat{\operatorname{vol}}(\overline{L})=\sup_{p\in\mathbb{N}}\widehat{\operatorname{vol}}(\overline{E}_{\bullet}^{(p)}).$
@@
In order to prove the assertion, we will prove that the restriction of $\nu_{\mathrm{sp}}$ on $[0,+\infty)$ can be approximated well by $\nu_{\mathrm{sp}}^{(p)}$ for $p\in\mathbb{N}$.
@@
be a division of the interval $[0,\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{E}_{\operatorname{HN},\bullet}))$ such that
@@
where the last equality is from [3, Chap. IV §5 $n^{\circ}$ 12 Proposition 22].
@@
$\max\{t,0\}-\max\{t-\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{E}_{\operatorname{HN},\bullet}),0\}$
when $j\to\infty$. By definition, the support of $\nu_{\mathrm{sp}}$ is bounded from above by $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{E}_{\operatorname{HN},\bullet})$, we obtain
@@
For a general graded linear series, if it satisfies the property of Proposition 4.2, we say that this graded linear series satisfies Fujita approximation. We would like to emphasize that this notion is different from those in [7, Definition 3.7] or [8, §6], though they are somewhat similar.
@@ 5 Fujita approximation of adelic line bundles
From the Fujita approximation of graded linear series proved in Proposition 4.2, we will prove the arithmetic Fujita approximation of an arithmetically big line bundle with coefficients in $\mathbb{Z}$, $\mathbb{Q}$ and $\mathbb{R}$.
@@ 5.1 Fujita approximation of big adelic line bundles
First, we will prove the arithmetic Fujita approximation of usual adelic line bundles. We will prove the nef version first, and apply it to prove the ample version. In order to apply Proposition 3.4.5, we assume the base field is perfect.
@@
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve with a perfect $K$, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. Let $X$ be a projective variety with $\dim(X)=d$, and $\overline{L}$ be an arithmetically big line bundle on $X$ over $S$. Then for every $\epsilon>0$, there exists a birational morphism $v:X^{\prime}\rightarrow X$ together with a positive integer $p$, such that the decomposition $v^{*}(\overline{L}^{\otimes p})=\overline{M}\otimes\overline{G}$ satisfies:
@@
We devide the proof into three steps.
@@
$X_{p}=\operatorname{Proj}\left(\operatorname{Im}\left(\bigoplus_{n\geqslant 0}\pi^{*}(E_{\operatorname{HN},np}^{(p)})\longrightarrow\bigoplus_{n\geqslant 0}L^{\otimes np}\right)\right).$
@@
We consider $\overline{E}^{(p)}_{\bullet}$ as a normed graded linear series over $S$, which is equipped with the subspace norms of $\overline{E}_{\bullet}=\bigoplus\limits_{D\geqslant 0}\overline{E}_{D}$. For any $n\geqslant 1$ and $w\in\Omega$, let .$\|\raisebox{1.72218pt}{.}\|_{n,w}$ be the quotient norm on $G_{p,w}=G_{p}\times_{\operatorname{Spec}K}\operatorname{Spec}K_{w}$ induced by the surjective homomorphism
@@
In fact, for any $x\in X_{p,w}(\overline{K}_{w})$ outside the exceptional divisor, it corresponds to a $1$-dimensional quotient of $E_{\operatorname{HN},p}^{0}$. For any $n\geqslant 1$, this quotient induces a $1$-dimensional quotient $\ell_{n,x}$ of $E_{\operatorname{HN},np}^{(p)}$.
By Hahn-Banach theorem, every linear form on $\ker(E_{\operatorname{HN},np}^{(p)}\rightarrow\ell_{n,x})$ can be extended to a linear form on $E_{\operatorname{HN},np}^{(p)}$. In other words, there exists a section $w\in E_{\operatorname{HN},np}^{(p)}$, whose image in $G_{p,w}^{\otimes n}(x)$ has norm $\|w\|_{\sup,w}$. As a section of $L_{w}^{\otimes np}$ on $X_{v}(\overline{K}_{w})$, we have $\|w(x)\|_{w}\leqslant\|w\|_{\sup,w}$, where .$\|\raisebox{1.72218pt}{.}\|_{w}$ is the metric on $\overline{L}$ indexed by $v\in\Omega$. Then for any section $u\in G_{p,w}$ over a neighborhood of $x$, we may assume that $u(x)^{\otimes n}$ equals to the image of $w(x)$ in $G_{p,w}^{\otimes n}(x)$ by dilation, then we have
@@
Let $\alpha_{n}=(\|\raisebox{1.72218pt}{.}\|_{n,w})_{w\in\Omega}$ be a family of adelic metrics on $G_{p}$, and we define $(M_{p},\beta_{n})=\phi_{p}^{*}(\overline{L})\otimes(G_{p},\alpha_{n})^{\vee}$. In this sense, $s$ is an effective section in $(M_{p},\beta_{n})$, then it is an effective adelic line bundle.
@@
$\widehat{\operatorname{vol}}(E_{\bullet}^{(p)},\alpha_{n})=\lim_{m\rightarrow\infty}\frac{\widehat{\deg}_{+}(E_{\operatorname{HN},mp}^{(p)},(\|\raisebox{1.72218pt}{.}\|_{n,\sup,w})_{w\in\Omega})}{m^{d+1}/(d+1)!},$
and .$\|\raisebox{1.72218pt}{.}\|_{n,\sup,w}$ is the superior norm induced by $\alpha_{n}=(\|\raisebox{1.72218pt}{.}\|_{n,w})_{w\in\Omega}$ on $G_{p}$. Then by [4, Corollary A.2], we have
@@
for infinitely many $n\geqslant N_{0}$.
In addition, since $\widehat{\mu}_{\min}(\overline{E}_{\operatorname{HN},np}^{(p)})\geqslant 0$ for all $n\geqslant 1$, then it has a positive asymptotic minimal slope. Then by Proposition 3.4.5, we obtain that $(G_{p},\alpha_{n})$ is arithmetically nef with the fact that $K$ is perfect.
@@
Next, we obtain the following ample version from Theorem 5.1 directly. The idea is absorbed from [4, Remark 4.4] and [11, Corollary 8.5.7].
@@
Let $\overline{B}$ be an arithmetically big line bundle on $X$. Then by [9, Corollary 6.4.10], for $m\in\mathbb{N}$ sufficiently large, we have $\overline{L}^{\otimes m}\otimes\overline{B}^{\vee}$ is arithmetically big. By Theorem 5.1, for any $\epsilon>0$, there exists a birational morphism $\phi:X^{\prime}\rightarrow X$, a $p\in\mathbb{N}$, and a decomposition
@@
By the continuity of the arithmetic volume function (see [9, Corollary 6.4.10]), we have the result. ∎
@@ 5.2 Case of $\mathbb{K}$-line bundles
In this part we will prove the arithmetic Fujita approximation is verified for adelic $\mathbb{Q}$-line bundles and $\mathbb{R}$-line bundles over adelic curves.
@@
First, we prove the case of adelic $\mathbb{Q}$-line bundle, which is a quite straightforward corollary from the usual case.
@@
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve with a perfect $K$, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. Let $X$ be a projective variety over $\operatorname{Spec}K$ with $\dim(X)=d$, and $\overline{L}$ be an arithmetically big $\mathbb{Q}$-line bundle on $X$ over $S$. For every $\epsilon>0$, there exists a birational morphism $\phi:X^{\prime}\rightarrow X$ together with a $p\in\mathbb{N}^{+}$, such that the decomposition $\phi^{*}(\overline{L}^{\otimes p})=\overline{M}\otimes\overline{G}$ satisfies:
@@
    $\overline{M}$ and $\overline{G}$ are adelic line bundles, where $\overline{M}$ is effective and $\overline{G}$ is arithmetically ample;
@@
By definition, there exists an $m\in\mathbb{N}$, such that $\overline{L}^{\otimes m}$ is a usual arithmetically big line bundle. Then we obtain the assertion by applying Corollary 5.1 to $\overline{L}^{\otimes m}$. ∎
@@
Next, we prove the case of adelic $\mathbb{R}$-line bundle, and see [9, §6.4.3] for an introduction to this notion. For this target, we prove an auxiliary lemma first, which allows us to approximate an arithmetically big $\mathbb{R}$-line bundle by an arithmetically big $\mathbb{Q}$-line bundle along an effective direction.
@@
The adelic $\mathbb{R}$-line bundle $\overline{L}$ is arithmetically big, so $L$ is a big $\mathbb{R}$-line bundle. By an equivalent definition [16, Definition 2.2.21], we have $L=\sum\limits_{i=1}^{r}a_{i}L_{i}$, where $a_{1},\ldots,a_{r}>0$ and $L_{1},\ldots,L_{r}$ are usual big line bundles. By definition, $L_{1},\ldots,L_{r}$ are all effective in the geometric sense.
We equip metrics on $L_{1},\ldots,L_{r}$ over $S$, such that $\overline{L}_{1},\ldots,\overline{L}_{r}$ are arithmetically effective. Then for any $b_{1},\ldots,b_{r}\in\mathbb{R}^{+}$, we have $b_{1}\overline{L}_{1}+\cdots+b_{r}\overline{L}_{r}$ is effective.
By the continuity of $\widehat{\operatorname{vol}}(\raisebox{1.72218pt}{.})$ from [9, Corollary 6.4.24], for all $\epsilon>0$, there exist $\delta_{1},\ldots,\delta_{r}>0$, such that for all $t_{i}\in(0,\delta_{i})$, $i=1,\ldots,r$, we always have
$0<\widehat{\operatorname{vol}}(\overline{L})-\widehat{\operatorname{vol}}\left(\overline{L}-\sum_{i=1}^{r}t_{i}\overline{L}_{i}\right)<\epsilon.$
If there does not exist any arithmetically big $\mathbb{Q}$-line bundles satisfying the requirements of the lemma, then $\overline{L}-\sum\limits_{i=1}^{r}t_{i}\overline{L}_{i}$ can never be a $\mathbb{Q}$-line bundle for all $t_{i}\in(0,\delta_{i})$, $i=1,\ldots,r$. It means there is no $\mathbb{Q}$-point in the region
$(a_{1}-\delta_{1},a_{1})\times\cdots\times(a_{r}-\delta_{r},a_{r})\subset\mathbb{R}^{r},$
which has a strict positive Lebesgue measure. Then it contradicts to the fact that $\mathbb{Q}^{r}$ is dense in $\mathbb{R}^{r}$ with respect to the Euclidean topology.
By choosing the constant $\epsilon$ sufficiently small, we pick a family $t_{1},\ldots,t_{r}\in\mathbb{R}^{+}$ such that $\overline{L}-\sum\limits_{i=1}^{r}t_{i}\overline{L}_{i}$ is an arithmetically big $\mathbb{Q}$-line bundle. By the above construction, $\sum\limits_{i=1}^{r}t_{i}\overline{L}_{i}$ is effective. Let
$\overline{L}^{\prime}=\overline{L}-\sum\limits_{i=1}^{r}t_{i}\overline{L}_{i},$
@@
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve with a perfect $K$, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. Let $X$ be a projective variety over $\operatorname{Spec}K$ with $\dim(X)=d$, and $\overline{L}$ be an arithmetically big adelic $\mathbb{R}$-line bundle on $X$ over $S$. For every $\epsilon>0$, there exists a birational morphism $\phi:X^{\prime}\rightarrow X$ together with a $p\in\mathbb{N}^{+}$, such that the decomposition $\phi^{*}(\overline{L}^{\otimes p})=\overline{M}\otimes\overline{G}$ satisfies:
@@
    $\overline{M}$ is an arithmetically effective $\mathbb{R}$-line bundle, and $\overline{G}$ is an arithmetically ample line bundle;